Solve the following equations:
-p/3 = 5 …...........

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

We have to multiply both sides of the equation by - 3,
Then, we get,
= - p/3 × (- 3) = 5 × (- 3)
= p = - 15

Thanks
Prachi Singh
2 years ago
Dear Student,
 
The following expression can be solved as follows :
-p/3 = 5
P = 5 × -3
P = -15
 
Hope it helps!
Thank you